x^2-4x+4y-8=0 standard form
Answer by ewatrrr(24785)

 
Hi,
x^2-4x+4y-8=0
(x-2)^2 - 4 +4y - 8 = 0
4y = -(x-2)^2 +12
y = -.25(x-2)^2 + 3